Find the sum of the solutions of the equation
(3x - 8/x)^2 - 7(3x - 8/x)^2 +10 = 0

Answer by stanbon(75887)

Find the sum of the solutions of the equation
(3x - 8/x)^2 - 7(3x - 8/x)^2 +10 = 0
----------------
Let w = (3x-8/x)
----------------
Substitute:
w^2 - 7w + 10 = 0
(w-5)(w-2) = 0
w = 5 or w = 2
------------------------
Solve for "x":
If w = 5
then 3x-8/x = 5
3x^2 - 5x - 8 = 0
(3x-8)(x+1) = 0
x = 8/3 or x = -1
--------------------------
If w = 2
then 3x-8/x = 2
3x^2 -2x -8 = 0
(x-2)(3x+4) = 0
x = 2 or x = -4/3
========================
Cheers,
Stan H.